OpenStudy (anonymous):

Please help! Indentify the rule for the function graphed below. y = x + 8 y = -x y = x - 8

OpenStudy (anonymous):

Here the graph.

OpenStudy (anonymous):

Every funtion pairs ___________ output value with one input value. more than one. at least one. exactly one.

OpenStudy (anonymous):

Which set of output values correctly complete the function table? -24, -8, -12 24, 8, 12 24, 8, -12

ganeshie8 (ganeshie8):

which q you need help wid

OpenStudy (anonymous):

#2

ganeshie8 (ganeshie8):

ok, by definition a function pairs pairs only one output value for each input value.

OpenStudy (anonymous):

Now can someone help me with #3

ganeshie8 (ganeshie8):

look at the table. there is a clue at the top of table : y=-4x

ganeshie8 (ganeshie8):

we need to use that to figure out the answer

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

is the x is -4 times?

ganeshie8 (ganeshie8):

yes, y is -4 times x

OpenStudy (anonymous):

cool

ganeshie8 (ganeshie8):

so, when x = -6, y = -4 x -6 = ?

OpenStudy (anonymous):

28

ganeshie8 (ganeshie8):

you mean 24 ?

OpenStudy (anonymous):

yea that

ganeshie8 (ganeshie8):

good :) so thats your output value for -6

ganeshie8 (ganeshie8):

look at table, whats the next x value ?

OpenStudy (anonymous):

8!

OpenStudy (anonymous):

thankyou

ganeshie8 (ganeshie8):

brilliant !

OpenStudy (anonymous):

Help! im stuck on #1

ganeshie8 (ganeshie8):

ok lets have a good look at the line

ganeshie8 (ganeshie8):

where the line is cutting y axis ?

ganeshie8 (ganeshie8):

y axis is that vertical thick line btw..

OpenStudy (anonymous):

the one crossing it is the one that lay down like a pencil the x axis.

ganeshie8 (ganeshie8):

yes, im asking at what point the red line is crossing the y axis

ganeshie8 (ganeshie8):

look at the graph, it is crossing at -8, right ?

OpenStudy (anonymous):

wait hold on

ganeshie8 (ganeshie8):

so thats the value when x = 0

ganeshie8 (ganeshie8):

and if you see, the red line is slowing rising up

OpenStudy (anonymous):

what was the question again?

ganeshie8 (ganeshie8):

we need find the equation for that red line

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

thats the part im stuck on

ganeshie8 (ganeshie8):

the red line is rising 1 grid up, for every 1 grid movement of right side.

ganeshie8 (ganeshie8):

so equation will be y = x - 8

ganeshie8 (ganeshie8):

see if that makes some sense

ganeshie8 (ganeshie8):

np

OpenStudy (anonymous):

The value of the dependent variable in a function is always ______ the value of the independent variable. determined by greater than equal to less than

OpenStudy (anonymous):

y=mx+b From the plot, the line crosses the y axis at -8. y=mx-8 Using the red dots on the plot, the slope, m, is 2/2 or 1. y=x-8